OpenStudy (anonymous):

find the exact value of the composition. arccos (cos pi/2)

OpenStudy (anonymous):

arccos(cos(pi radians) / 2) = 2.0943951 rad

OpenStudy (anonymous):

π 0 pi divided by three pi divided by two

OpenStudy (anonymous):

isnt in my answer choice

OpenStudy (anonymous):

@Tazr

OpenStudy (anonymous):

what do you get when you compose a function with its inverse?

OpenStudy (anonymous):

u get x

OpenStudy (ybarrap):

What is x in your case?

OpenStudy (ybarrap):

You're on the right track: f(g(x))=x, here f=acos(x), g=cos(x)

OpenStudy (anonymous):

is it pi/2

OpenStudy (ybarrap):

You already said it earlier, for inverses, you get x back Nice job.